OpenStudy (anonymous):

h(x) = −3x − 10, find h(−3). −19 −1 1 2.3

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

substitue the value of x to -3x -10 h(-3) = -3(-3)-10 h(-3) = 9-10 h(-3) = -1

OpenStudy (anonymous):

If N = {integers} and subset C = {integers divisible by 10}, what is C'? {integers except those divisible by 10} { } {10, 20, 30, 40, …} {integers}